# FLoRA: An Advanced AI-Powered Engine to Facilitate Hybrid Human-AI Regulated Learning

### v2

@@ Abstract
SRL, defined as learners’ ability to systematically plan, monitor, and regulate their learning activities, is crucial for sustained academic achievement and lifelong learning competencies. Emerging Artificial Intelligence (AI) developments profoundly influence SRL interactions by potentially either diminishing or strengthening learners’ opportunities to exercise their own regulatory skills. Recent literature emphasizes a balanced approach termed Hybrid Human-AI Regulated Learning (HHAIRL), in which AI provides targeted, timely scaffolding while preserving the learners’ role as active decision-makers and reflective monitors of their learning process. Nevertheless, existing digital tools frequently fall short, lacking adaptability, focusing narrowly on isolated SRL phases, and insufficiently support meaningful human-AI interactions. In response, this paper introduces the enhanced FLoRA Engine, which incorporates advanced Generative Artificial Intelligence (GenAI) features and state-of-the-art learning analytics, explicitly grounded in SRL and HHAIRL theories. The FLoRA Engine offers instrumentation tools such as collaborative writing, multi-agents chatbot, and detailed learning trace logging to support dynamic, adaptive scaffolding tailored to individual needs in real time. We further present a summary of several research studies that provide the validations for and illustrate how these instrumentation tools can be utilized in real-world educational and experimental contexts. These studies demonstrate the effectiveness of FLoRA Engine in fostering SRL and HHAIRL, providing both theoretical insights and practical solutions for the future of AI-enhanced learning context.
@@ 1 Introduction
Self-regulated learning (SRL) is often defined as a learner’s ability to plan, monitor, and adjust personal learning processes in order to meet self-determined goals (Azevedo and Wiedbusch, 2023). It encompasses an array of cognitive and metacognitive strategies – such as goal setting, strategic decision-making, and self-monitoring – that together help learners direct their own progress (Alvarez et al., 2022). Through the active exercise of these strategies, learners can deepen their understanding, build resilience, and enhance their capacity to transfer acquired knowledge and skills across contexts (Winne, 2017b). Moreover, SRL is not only pivotal for achieving higher academic performance – extensive research indicates that well-regulated learners generally achieve better outcomes (Dent and Koenka, 2016; Xu et al., 2023; Broadbent and Poon, 2015) – but also for lifelong learning, which has gained heightened importance in an era where constant skill adaptation is a necessity (Seufert, 2018).
@@
Generative Artificial Intelligence (GenAI) and other recent AI advances are reshaping how educators, researchers, and learners view SRL (Järvelä et al., 2023). On one hand, many adaptive learning technologies can offload tasks like monitoring and selecting new materials, which can inadvertently reduce learners’ opportunities to exercise self-regulation (Molenaar et al., 2023b). On the other hand, the integration of data-driven AI support with explicit scaffolds aimed at promoting and reinforcing learners’ own regulatory skills (Molenaar, 2022a; Lim et al., 2024). The AI can step in with timely, data-driven insights while still encouraging learners to make key decisions and reflect on their learning process (Molenaar and Wise, 2022). This synergy leverages the strengths of both AI and the human learner, ensuring that the learner remains at the centre and continues to build the metacognitive monitoring and control skills essential for future learning, with the process termed as Hybrid Human-AI Regulated Learning (HHAIRL) (Huang, 2024; Nguyen, 2025; Järvelä et al., 2023; Molenaar and Wise, 2022). Still, there is a risk that an excessive reliance on AI could diminish learners’ engagement in important self-regulatory activities (Zhang et al., 2024), such as monitoring their own progress or planning their learning, thereby impeding the growth of robust SRL skills (Authors, 2024a).
@@
Despite increasing attention to Hybrid Human-AI Regulation (Huang, 2024; Nguyen, 2025; Järvelä et al., 2023; Molenaar and Wise, 2022), several gaps and challenges remain. First, existing technologies for the real-time identification of a learner’s SRL are relatively underdeveloped. Many current tools lack adaptability; they often work well for specific tasks but do not account for the diverse realities of educational environments, where students engage with varied materials and may frequently learn collaboratively (Alvarez et al., 2022). In addition, many tools focus on particular SRL phases rather than encompassing the entire cycle of setting goals, monitoring progress, and evaluating outcomes (Alvarez et al., 2022). The design of tools like dashboards that help learners understand and exert control alongside the AI is still at an early stage; simply showing performance metrics is not enough (Matcha et al., 2020). Furthermore, the emergence of AI especially GenAI, capable of providing human-like interactions and demonstrations of expert-level reasoning, significantly alters typical human-tool interaction patterns (Edwards et al., 2025; Kasneci et al., 2023). Learners increasingly trust and rely on GenAI systems due to their sophisticated capacities and human-like interactions, fundamentally changing how learners engage in SRL processes such as goal-setting, monitoring, and reflection (Hauske and Bendel, 2024; Lai, 2024). Nonetheless, existing SRL support tools, primarily designed prior to these advancements, are insufficiently equipped with GenAI-based functionalities to reliably track, interpret, and scaffold these complex human-GenAI interactions (Edwards et al., 2025; Holmes et al., 2023). Finally, many SRL support tools often exhibit strong dependencies on their specific learning contexts (Wong et al., 2019; Matcha et al., 2020), limiting their applicability across diverse disciplines and making them unsuitable for evaluating the effectiveness of HHAIRL in diversified learning scenarios (Molenaar, 2022a).
@@
In response to these limitations, the FLoRA Engine has been developed based on the previous work (Authors, 2025a). Building upon prior version, the current version of FLoRA not only significantly expands the three main modules – instrumentation tools, trace parser, and scaffolding, but also introduces several innovative features to support HHAIRL research. One particularly significant advancement in this iteration is the integration of GenAI technologies, such as GPT-4o, which enable more dynamic, personalised, and adaptive learner-centred support. In parallel, FLoRA incorporates advanced learning analytics methodologies (Gašević et al., 2015; Matcha et al., 2019) grounded firmly in established SRL theories (Saint et al., 2020b; Siadaty et al., 2016), ensuring that insights generated from learner-interaction data remain theoretically coherent and practically meaningful. By aligning closely with emerging trends in AI-enhanced educational practice, FLoRA serves a dual purpose: on one hand, it provides researchers with a comprehensive, robust platform for systematically investigating the complex dynamics of Hybrid Human-AI regulation; on the other, it offers educators and learning scientists a practical example illustrating how GenAI technologies can effectively enhance SRL-supportive educational contexts.
@@ 2.1 Self-Regulated Learning
The importance of SRL lies in its significant impact on educational outcomes and lifelong learning (Dent and Koenka, 2016; Xu et al., 2023; Broadbent and Poon, 2015; Guo, 2022). SRL equips learners with the skills necessary to adapt to various learning contexts and challenges, fostering autonomy and resilience. When students practice SRL, they are more likely to exhibit higher levels of self-efficacy, persistence, and academic achievement (Uzir et al., 2020; van der Graaf et al., 2022). Moreover, SRL promotes deeper engagement with learning materials (Li and Lajoie, 2022), critical thinking (Anwar and Muti’ah, 2022), and the ability to transfer knowledge to new situations (Stebner et al., 2022; Schuster et al., 2020). As educational environments become increasingly complex, the ability to self-regulate is essential for success. Educators recognise that teaching students how to learn, rather than just what to learn, empowers them to become effective, lifelong learners who can navigate and thrive in a rapidly changing world (Guo, 2022).
@@ 2.2 The Detection and Measurement of SRL
One effective approach to detecting SRL is the utilization of log data and learning analytics within digital learning platforms (Siadaty et al., 2016; Winne, 2022; Bernacki, 2025). As learners interact with computer-based educational resources, they generate extensive trace data—clickstreams, navigation paths, and tool usage logs—that can be analysed to infer SRL activities (Papamitsiou and Economides, 2014; Bannert et al., 2014; Azevedo and Gašević, 2019; Li et al., 2020; Deininger et al., 2025). For instance, if a learner chooses to attempt a quiz before engaging with reading materials, this might indicate proactive planning and goal-setting behaviours. However, not all SRL actions leave discernible traces in the log data, and some SRL processes, particularly metacognitive ones, may remain hidden or indistinct, making it challenging to achieve a comprehensive understanding solely through raw interaction data (Winne, 2017a, 2018).
@@
Analysing the rich trace data generated requires sophisticated methods to accurately map observable behaviours to underlying SRL processes (Bannert, 2007; Molenaar et al., 2023a). Conceptualizing SRL as a series of events with increasing complexity—occurrence, contingency, and patterned contingency—provides a structured framework for this analysis (Lim et al., 2021). The occurrence level captures the frequency of specific learning actions, such as the number of times a learner highlights text. The contingency level examines sequences of actions, providing context by looking at how one action follows another—for example, highlighting text immediately after taking a quiz. The patterned contingency level involves identifying recurring sequences or patterns of actions that suggest a strategic approach to learning, such as a consistent cycle of reading, note-taking, and self-testing. By employing temporal analytical methods like process mining and Markov models, researchers can uncover these patterns within the data, revealing how learners self-regulate over time (Saint et al., 2022).
@@ 2.3 Hybrid Human-AI and Socially Shared Regulated Learning
Socially-Shared Regulation of Learning (SSRL) is the process through which members of a group collectively plan, monitor and adapt their activity in order to reach a shared learning goal (Järvelä et al., 2018; Haataja et al., 2022). Regulation decisions—clarifying the task, setting sub-goals, checking mutual understanding or reallocating effort—are negotiated and enacted together, not left to any single individual (Zheng et al., 2019; Kent and Cukurova, 2020). Traditionally, teachers have scaffolded these collective metacognitive moves; now GenAI can join the conversation (Järvelä et al., 2023). By analysing multimodal data streams in real time, an AI agent can detect moments when the group drifts apart, highlight emerging misconceptions or surface silent contributions, effectively acting as an additional “group member” that nudges the team toward productive dialogue (Edwards et al., 2025). Thus, AI becomes an enabler of richer SSRL, augmenting human awareness rather than replacing it.
@@
Attending to SSRL and HHAIR is therefore crucial in contemporary education. Research shows that learners often under-utilise effective regulation strategies, while many adaptive tools silently take those decisions away, depriving students of practice (Nazaretsky et al., 2024). Thoughtfully designed AI tools and systems need to address both problems: it keeps collaboration efficient in the moment, provides personalised and timely feedback, and explicitly models the metacognitive moves for learners to internalise (Cukurova, 2024; Yan et al., 2024). However, these tools remains limited. One key reason for this gap is the complexity of designing adaptive AI systems that can dynamically share control with human learners and educators (Hao and Cukurova, 2023). Current AI-driven educational technologies are heavily focused on providing automated and fixed recommendations, often lacking mechanisms for reciprocal interaction and real-time feedback integration from learners (Roll and Winne, 2015). Most existing adaptive learning systems prioritise knowledge acquisition and task adaptation, rather than supporting SRL and metacognitive processes, which require the AI to recognise, interpret, and respond to a wide range of learner behaviours, emotions, and cognitive states dynamically (Azevedo and Gašević, 2019). Developing systems that can contextually adjust the level of intervention based on individual’s skills and autonomy remains a major technical challenge. Furthermore, most learning management systems (LMS) and AI-driven tutoring platforms lack transparent, explainable AI mechanisms to ensure that learners and teachers can understand how AI-driven decisions are made (Shin, 2021; Khosravi et al., 2022). Without clear communication of AI reasoning, users may struggle to trust or effectively engage with the system (Kizilcec, 2016). Limited interdisciplinary collaboration among AI developers, cognitive scientists, and education practitioners further contributes to the slow development of robust hybrid regulation tools (Molenaar, 2021).
@@ 3.1 From Learning Analytics to Personalised Scaffolds with GenAI
FLoRA Engine incorporates an analytics-based GenAI scaffolding function that integrates learning analytics and GenAI to offer personalised support for students’ SRL processes. The tool is shown in Figure 2. Its name has been updated to "Instruction Panel" to enhance user clarity. At predefined intervals, users receive scaffolding messages through this panel, providing personalized guidance during their learning process.
The mechanism that triggers this support is illustrated in Figure 3. As learners work on the given tasks, learning analytics techniques collect real-time data, capturing both internal learning conditions and SRL processes. According to Winne (2022), internal conditions comprise, for instance, a learner’s level of domain knowledge (a static condition) and their awareness of time constraints (a dynamic condition). In a time-limited learning task, a learner might initially adopt a linear approach, but as they become aware of the remaining time, they may adjust their strategies to accommodate this constraint. Consequently, time-awareness can be considered a dynamic learning condition that updates over the course of the task. In FLoRA , such changes in dynamic conditions are detected in real-time through trace data (e.g., a click on the timer to check the remaining time). Meanwhile, static conditions, such as learners’ domain knowledge levels, are determined through pre-task surveys and assessments. FLoRA automatically scores these instruments, and the resulting scores feed into the GenAI-based scaffolding tool to tailor the support provided to each learner. Each learning condition is converted into a predefined statement and added to the GenAI prompt. More details can be found in Appendix C. For example, if a student’s trace data shows they clicked on the timer, it indicates awareness of time constraints, generating “the student is aware of the time constraint.” Conversely, no click indicates unawareness, resulting in “the student is not aware of the time constraint.”
@@
Meanwhile, SRL processes are captured in real-time representing how learners self-regulate their learning. For example, when a student working on drafting their essay, and checking task instruction and rubrics during the drafting, this sequential process will be automatically labelled as evaluation. The current FLoRA is compatible with different theoretical models informing the design of trace parser, including Winne and Hadwin’s COPES model (2022) and Zimmerman’s cyclical model of SRL (2013), which is empirically validated using think-aloud data as the reference points (Authors, 2021). Our previous lab study identified critical SRL processes that were significantly associated with learning performance at various stages of the task (Authors, 2024b). For instance, in a timed reflective writing task, students who engaged in orientation—using task instructions and marking rubrics to guide their reading within the first two minutes—scored significantly higher than those who did not (van der Graaf et al., 2023). Drawing on these findings, FLoRA can automatically trigger GenAI-powered scaffolding whenever these critical SRL processes remain undetected for a specified time period.
@@
As shown in Figure 3, learning conditions and critical SRL processes—collected through real-time analytics—are sent to the FLoRA system, which automatically generates a GenAI prompt using the relevant template. These prompts are then passed to GenAI to produce responses that become SRL scaffolding. The resulting feedback is delivered to students via the GenAI-based scaffolding tool, nudging them toward effective SRL processes with content tailored to their evolving needs (Figure 2). This integrated design—combining learning analytics with GenAI—offers clear advantages over rule-based scaffolding, which relies on a predetermined set of messages (Srivastava et al., 2022; van der Graaf et al., 2023). Since rule-based systems require all potential messages to be predefined, addressing multiple learning conditions simultaneously becomes unwieldy and impractical due to the sheer complexity of enumerating possible rule combinations (Backhaus et al., 2017).

### v3

@@ Abstract
Self-Regulated Learning (SRL), defined as learners’ ability to systematically plan, monitor, and regulate their learning activities, is crucial for sustained academic achievement and lifelong learning competencies. Emerging AI developments profoundly influence SRL interactions by potentially either diminishing or strengthening learners’ opportunities to exercise their own regulatory skills. Recent literature emphasizes a balanced approach termed Hybrid Human-AI Regulated Learning (HHAIRL), in which AI provides targeted, timely scaffolding while preserving the learners’ role as active decision-makers and reflective monitors of their learning process. Central to HHAIRL is the integration of adaptive and personalized learning systems; by modelling each learner’s knowledge and self-regulation patterns, AI can deliver contextually relevant scaffolds that support learners during all phases of the SRL process. Nevertheless, existing digital tools frequently fall short, lacking adaptability and personalisation, focusing narrowly on isolated SRL phases, and insufficiently supporting meaningful human-AI interactions. In response, this paper introduces the enhanced FLoRA Engine, which incorporates advanced generative AI features and state-of-the-art learning analytics, and grounds in solid educational theories. The FLoRA Engine offers tools such as collaborative writing, multi-agent chatbots, and detailed learning trace logging to support dynamic, adaptive scaffolding of self-regulation tailored to individual needs in real time. We further present a summary of several research studies that provide the validations for and illustrate how these tools can be utilized in real-world educational and experimental contexts. These studies demonstrate the effectiveness of FLoRA Engine in fostering SRL, providing both theoretical insights and practical solutions for the future of AI-enhanced learning contexts.
@@ 1 Introduction
Self-regulated learning (SRL) is the ability to plan, monitor, and adjust one’s learning by coordinating cognitive, metacognitive, motivational, and emotional processes (Azevedo and Wiedbusch, 2023; Du et al., 2023). Learners with strong SRL set clear goals, track their progress, and adapt strategies, which typically leads to better outcomes across tasks and over time (Xu et al., 2023; Zhao et al., 2025). Given the importance of SRL and the rapid evolution of digital learning environments and tools, it is crucial to reconsider how SRL is understood and supported (Seufert, 2018; Winne, 2017b). In this context, advances in Generative Artificial Intelligence (GenAI) are particularly significant, as they are beginning to reshape how educators, researchers, and learners conceptualize and foster SRL (Järvelä et al., 2023). On the one hand, many adaptive learning technologies can offload SRL processes like monitoring and selecting new materials, which can inadvertently reduce learners’ opportunities to exercise self-regulation (Molenaar et al., 2023b). On the other hand, the integration of data-driven AI support with explicit scaffolds aimed at promoting and reinforcing learners’ own regulatory skills (Molenaar, 2022a; Lim et al., 2024). The AI can step in with timely, data-driven insights while still encouraging learners to make key decisions and reflect on their learning process (Molenaar and Wise, 2022). This synergy leverages the strengths of both AI and the human learner, ensuring that the learner remains at the centre and continues to build the metacognitive monitoring and control skills essential for future learning, with the process termed as Hybrid Human-AI Regulated Learning (HHAIRL) (Huang, 2024; Nguyen, 2025; Järvelä et al., 2023; Molenaar and Wise, 2022). Still, there is a risk that an excessive reliance on AI could diminish learners’ engagement in important self-regulatory activities (Zhang et al., 2024), such as monitoring their own progress or planning their learning, thereby impeding the growth of robust SRL skills (Fan et al., 2025).
@@
Despite increasing attention to Hybrid Human-AI Regulation (Huang, 2024; Nguyen, 2025; Järvelä et al., 2023; Molenaar and Wise, 2022), several gaps and challenges remain. First, existing technologies for the real-time identification of a learner’s SRL are relatively underdeveloped. Many current tools lack adaptability; they often work well for specific tasks but do not account for the diverse realities of educational environments, where students engage with varied materials and may frequently learn collaboratively (Alvarez et al., 2022). In addition, many tools focus on particular SRL phases rather than encompassing the entire cycle of setting goals, monitoring progress, and evaluating outcomes (Alvarez et al., 2022; Guan et al., 2025). The design of tools like dashboards that help learners understand and exert control alongside the AI is still at an early stage; simply showing performance metrics is not enough (Matcha et al., 2020). Furthermore, the emergence of AI especially GenAI, capable of providing human-like interactions and demonstrations of expert-level reasoning, significantly alters typical human-tool interaction patterns (Edwards et al., 2025; Kasneci et al., 2023). Learners increasingly trust and rely on GenAI systems due to their sophisticated capacities and human-like interactions, fundamentally changing how learners engage in SRL processes such as goal-setting, monitoring, and reflection (Hauske and Bendel, 2024; Lai, 2024). Nonetheless, existing SRL support tools, primarily designed prior to these advancements, are insufficiently equipped with GenAI-based functionalities to reliably track, interpret, and scaffold these complex human-GenAI interactions (Edwards et al., 2025; Holmes et al., 2023; Tsai et al., 2025). Finally, many SRL support tools often exhibit strong dependencies on their specific learning contexts (Wong et al., 2019; Matcha et al., 2020; Geng and Su, 2025), limiting their applicability across diverse disciplines and making them unsuitable for evaluating the effectiveness of HHAIRL in diversified learning scenarios (Molenaar, 2022a).
@@
In response to these limitations, the FLoRA Engine has been developed based on the previous work (Li et al., 2025c). Building upon prior version, the current version of FLoRA not only significantly expands the three main modules – instrumentation tools, trace parser, and scaffolding, but also introduces several innovative features to support HHAIRL research. One particularly significant advancement in this iteration is the integration of GenAI technologies, such as GPT-4o, which enable more dynamic, personalised, and adaptive learner-centred support. In parallel, FLoRA incorporates advanced learning analytics methodologies (Gašević et al., 2015; Matcha et al., 2019) grounded firmly in established SRL theories (Saint et al., 2020b; Siadaty et al., 2016), ensuring that insights generated from learner-interaction data remain theoretically coherent and practically meaningful. By aligning closely with emerging trends in AI-enhanced educational practice, FLoRA serves a dual purpose: on one hand, it provides researchers with a comprehensive, robust platform for systematically investigating the complex dynamics of Hybrid Human-AI regulation; on the other, it offers educators and learning scientists a practical example illustrating how GenAI technologies can effectively enhance SRL-supportive educational contexts.
@@ 2.1 Self-Regulated Learning
The importance of SRL lies in its significant impact on educational outcomes and lifelong learning (Dent and Koenka, 2016; Xu et al., 2023; Broadbent and Poon, 2015; Guo, 2022). SRL equips learners with the skills necessary to adapt to various learning contexts and challenges, fostering autonomy and resilience. When students practice SRL, they are more likely to exhibit higher levels of self-efficacy, persistence, and academic achievement (Uzir et al., 2020; van der Graaf et al., 2022). Moreover, SRL promotes deeper engagement with learning materials (Li and Lajoie, 2022), critical thinking (Anwar and Muti’ah, 2022), and the ability to transfer knowledge to new situations (Stebner et al., 2022; Schuster et al., 2020). As educational environments become increasingly complex, the ability to self-regulate is essential for learners to succeed in learning performance. Educators recognise that teaching students how to learn, rather than just what to learn, empowers them to become effective, lifelong learners who can navigate and thrive in a rapidly changing world (Guo, 2022).
@@ 2.2 The Detection and Measurement of SRL
One effective approach to detecting SRL is the utilization of log data and learning analytics within digital learning platforms (Siadaty et al., 2016; Winne, 2022; Bernacki, 2025). As learners interact with computer-based educational resources, they generate extensive trace data—clickstreams, navigation paths, and tool usage logs—that can be analysed to infer SRL activities (Papamitsiou and Economides, 2014; Bannert et al., 2014; Azevedo and Gašević, 2019; Li et al., 2020; Deininger et al., 2025). For instance, if a learner chooses to attempt a quiz before engaging with reading materials, this might indicate proactive planning and goal-setting behaviours. However, not all SRL actions leave discernible traces in the log data, and some SRL processes, particularly metacognitive ones, may remain hidden or indistinct, making it challenging to achieve a comprehensive understanding solely through raw interaction data (Winne, 2017a, 2018; Geng and Su, 2025).
@@
Analysing the rich trace data generated requires sophisticated methods to accurately map observable behaviours to underlying SRL processes (Bannert, 2007; Molenaar et al., 2023a). Conceptualizing SRL as a series of events with increasing complexity—occurrence, contingency, and patterned contingency—provides a structured framework for this analysis (Lim et al., 2021). The occurrence level captures the frequency of specific learning actions, such as the number of times a learner highlights text. The contingency level examines sequences of actions, providing context by looking at how one action follows another—for example, highlighting text immediately after taking a quiz. The patterned contingency level involves identifying recurring sequences or patterns of actions that suggest a strategic approach to learning, such as a consistent cycle of reading, note-taking, and self-testing. By employing temporal analytical methods like process mining and Markov models, researchers can uncover these patterns within the data, revealing how learners self-regulate over time (Saint et al., 2022; Li et al., 2025b).
@@ 2.3 Hybrid Human-AI and Socially Shared Regulated Learning
Socially-Shared Regulation of Learning (SSRL) is the process through which members of a group collectively plan, monitor and adapt their activity in order to reach a shared learning goal (Järvelä et al., 2018; Haataja et al., 2022). Regulation decisions—clarifying the task, setting sub-goals, checking mutual understanding or reallocating effort—are negotiated and enacted together, not left to any single individual (Zheng et al., 2019; Kent and Cukurova, 2020). Traditionally, teachers have scaffolded these collective metacognitive moves; now GenAI can join the conversation (Järvelä et al., 2023; Yan et al., 2025; Tsai et al., 2025). By analysing multimodal data streams in real time, an AI agent can detect moments when the group drifts apart, highlight emerging misconceptions or surface silent contributions, effectively acting as an additional “group member” that nudges the team toward productive dialogue (Edwards et al., 2025). Thus, AI becomes an enabler of richer SSRL, augmenting human awareness rather than replacing it.
@@
Attending to SSRL and HHAIR is therefore crucial in contemporary education. Research shows that learners often under-utilise effective regulation strategies, while many adaptive tools silently take those decisions away, depriving students of practice (Nazaretsky et al., 2024). Thoughtfully designed AI tools and systems need to address both problems: it keeps collaboration efficient in the moment, provides personalised and timely feedback, and explicitly models the metacognitive moves for learners to internalise (Cukurova, 2024; Yan et al., 2024). However, these tools remains limited. One key reason for this gap is the complexity of designing adaptive AI systems that can dynamically share control with human learners and educators (Hao and Cukurova, 2023). Current AI-driven educational technologies are heavily focused on providing automated and fixed recommendations, often lacking mechanisms for reciprocal interaction and real-time feedback integration from learners (Roll and Winne, 2015). Most existing adaptive learning systems prioritise knowledge acquisition and task adaptation, rather than supporting SRL and metacognitive processes, which require the AI to recognise, interpret, and respond to a wide range of learner behaviours, emotions, and cognitive states dynamically (Azevedo and Gašević, 2019). Developing systems that can contextually adjust the level of intervention based on individual’s skills and autonomy remains a major technical challenge. This intervention means the support offered by the systems such as feedback of learners’ current learning progress or scaffolding that can lead to cognitive and metacognitive learning process adjustments (Saint et al., 2024). Furthermore, most learning management systems (LMS) and AI-driven tutoring platforms lack transparent, explainable AI mechanisms to ensure that learners and teachers can understand how AI-driven decisions are made (Shin, 2021; Khosravi et al., 2022). Without clear communication of AI reasoning, users may struggle to trust or effectively engage with the system (Kizilcec, 2016). Limited interdisciplinary collaboration among AI developers, cognitive scientists, and education practitioners further contributes to the slow development of robust hybrid regulation tools (Molenaar, 2021).
While AI holds immense promise in education—such as supporting SSRL and HHAIR with personalised feedback and guidance—recent studies caution that heavy reliance on AI may lead to metacognitive disengagement, where learners offload reflective thinking to the AI. For example, Zou et al. (2024) found that students were more engaged and made more accurate revisions when using human teacher feedback compared to feedback from ChatGPT, suggesting that AI assistance, while helpful, might not stimulate the same level of active effort. Potential risks of intensive AI use in learning include a loss of student agency where learners may become passive and allow AI to make decisions for them (Lepage and Collin, 2024); the excessive automation of metacognitive processes which can prevent learners from practising self-monitoring or critical thinking because AI takes over these functions (Gerlich, 2025; Zhai et al., 2024); and the reproduction of biases, as AI models may mirror or even amplify existing societal biases in their feedback or suggestions, potentially leading to unfair or skewed learning experiences (Chinta et al., 2024; Suchithra and Arya, 2025).
@@ 3.1 From Learning Analytics to Personalised Scaffolds with GenAI
FLoRA Engine incorporates an analytics-based GenAI scaffolding function that integrates learning analytics and GenAI to offer personalised support for students’ SRL processes. The tool is shown in Figure 2. Its name has been updated to "Instruction Panel" to enhance user clarity. At predefined intervals, users receive scaffolding messages through this panel, providing personalised guidance during their learning process.
@@
The mechanism that triggers this support is illustrated in Figure 3. As learners work on the given tasks, learning analytics techniques collect real-time data, capturing both internal learning conditions and SRL processes. According to Winne (2022), internal conditions comprise, for instance, a learner’s level of domain knowledge (a static condition) and their awareness of time constraints (a dynamic condition). In a time-limited learning task, a learner might initially adopt a linear approach, but as they become aware of the remaining time, they may adjust their strategies to accommodate this constraint. Consequently, time-awareness can be considered a dynamic learning condition that updates over the course of the task. In FLoRA , such changes in dynamic conditions are detected in real-time through trace data (e.g., a click on the timer to check the remaining time). Meanwhile, static conditions, such as learners’ domain knowledge levels, are determined through pre-task surveys and assessments. FLoRA automatically scores these instruments, and the resulting scores feed into the GenAI-based scaffolding tool to tailor the support provided to each learner. Each learning condition (defined in Table 2) is converted into a predefined statement and added to the GenAI prompt. More details can be found in Appendix A.2. For example, if a student’s trace data shows they clicked on the timer, it indicates awareness of time constraints, generating “the student is aware of the time constraint.” Conversely, no click indicates unawareness, resulting in “the student is not aware of the time constraint.”
@@
SRL processes are continuously recorded in real time, providing insights into how learners actively manage and regulate their learning activities. The current version of FLoRA is designed to be compatible with various theoretical frameworks that inform the development of the trace parser. Notably, this includes Winne and Hadwin’s COPES model (2022), as well as Bannert’s framework (2007), which builds upon Zimmerman’s cyclical model of SRL (2013). To illustrate, consider an example based on Bannert’s framework: when a student is engaged in drafting an essay, and subsequently reviews the task instructions and rubrics before returning to continue writing, the raw interaction trace records these learning activities as WRITE_ESSAY and READ_INSTRUCTION. The sequential pattern WRITE_ESSAY → READ_INSTRUCTION → WRITE_ESSAY is then annotated as representing the Metacognition-Evaluation phase of the SRL process. A comprehensive explanation of how raw trace data are mapped onto specific SRL processes is provided by Li et al. (2025c), and this mapping procedure has been empirically validated using think-aloud protocols (Fan et al., 2022). Further, a recent laboratory study demonstrated that certain SRL processes are significantly correlated with improved learning outcomes at specific stages of task completion (Li et al., 2024). For example, in a timed reflective writing task, students who engaged in orientation behaviours—such as consulting task instructions and assessment rubrics within the first two minutes—achieved significantly higher scores than those who did not (van der Graaf et al., 2023). Building upon these findings, FLoRA has been designed to automatically trigger GenAI-powered scaffolding whenever critical SRL processes are not detected within a predefined time window, thereby providing timely and targeted support to enhance learning. The thresholds that trigger scaffolding are not fixed but are configurable and defined according to the specific design and aims of each study.
Figure 4: GenAI Prompt Example
As shown in Figure 3, learning conditions and critical SRL processes—collected through real-time analytics—are sent to the FLoRA system, which automatically generates a GenAI prompt using the relevant template. An example of a GenAI prompt is presented in Figure 4. In this template, the "Learning Conditions and SRL Process" section is dynamically generated, while the remaining components remain consistent across all instances. These prompts are then passed to GenAI to produce responses that become SRL scaffolding. The resulting feedback is delivered to students via the GenAI-based scaffolding tool, nudging them toward effective SRL processes with content tailored to their evolving needs (Figure 2). This integrated design—combining learning analytics with GenAI—offers clear advantages over rule-based scaffolding, which relies on a predetermined set of messages (Srivastava et al., 2022; van der Graaf et al., 2023). Since rule-based systems require all potential messages to be predefined, addressing multiple learning conditions simultaneously becomes unwieldy and impractical due to the sheer complexity of enumerating possible rule combinations (Backhaus et al., 2017).